Let A = 1 -2 31 2 1 -k . Find the value of k for which det A = 10. Your value of k should be an integer. Lo –3 kJ Answer:
Answer #1

I 1 2 3 La B Lok ] A1= 1.-2 at 0-3 kl 1(K-3R).+2 (2KJ+ 3(-6)=10 -2K 64 K - 18 = 10 2 K = 28 K= 14 1